17. A tenfold serial dilution is performed with a final dilution of 1/10000. The beginning dilution is the series in tube 1 is a 1/10 dilution. If the original concentration had been 25000 ng/dL, what would the concentration be in tube 3?
1
Expert's answer
2020-04-18T06:45:31-0400

1 dL =0.1 litres

original concentration=25000ng/dL = 25000"\times"10-9/0.1 g/L =2.5"\times"10-4 g/L

in the first tube it is diluted to 1/10 of the original hence the concentration will become

"\frac{1}{10}\times2.5\times10^{-4}=2.5\times10^{-5}g/L"

similarly in the 2nd second tube it would 1/10 of the previous concentration which will be

"\frac{1}{10}\times2.5\times10^{-5}=2.5\times10^{-6}g/L"

similarly in the 3rd second tube it would 1/10 of the previous concentration which will be

"\frac{1}{10}\times2.5\times10^{-6}=2.5\times10^{-7}g/L"